Anton is 4 times as old as patrick. in 10 years' time Anton will be twice as old as patrick. what is Anton's age?
Answer by ikleyn(52790)

.

x - Patrick's age

4x Anton's age


In 10 years


    2*(x + 10) = 4x+10


Simplify and solve


    2x + 20 = 4x + 10

    20 - 10 = 4x - 2x

    10      = 2x

     x      = 10/2 = 5.


ANSWER.  Patrick is 5 years old.  Anton is 20 years old.

Solved.

---------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.